# On estimation and order selection for multivariate extremes via clustering

### v3

@@ 6.2.2 River Discharge Data
The river discharge data concerns the daily discharge rate of rivers in North America sourced from the Global Runoff Data Centre [13]. The dataset comprises 16,386 daily records of discharge values from 13 stations spanning the period from December 1, 1976, to October 11, 2021. These 13 stations, shown in Table 1 and Figure 10, are positioned along 5 rivers in America: Willamette River, Mississippi River, Williamson River, Hudson River, and Broad River.
Table 1: River Discharge Stations
@@
Fig. 10: 13 River Discharge Stations
Fig. 11: Penalized ASW Curves for River Discharge Data (top $10\%$ norms)
Fig. 12: Estimated $B^{\top}$ for River Discharge Data (top $10\%$ norms)
As in the previous example, Figure 12 presents the penalized ASW curves, from which we found that 6 seems to be an appropriate choice of order. Figure 12 illustrates the cluster centers obtained from the $k$-pc algorithm when the order is chosen as $6$. In Figure 12 , we convert the spectral estimation to the factor matrix $B$ following the method in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$. In addition, for each row of the matrix $B$, we find to which factor index (the same as the cluster index in Figure 12) the largest value corresponds. We include these factor indices in the last column of Table 1, which can be viewed roughly as markings of groups of concomitant extremes. The results align well with the expected geographical context: These 6 groups are in good accordance with the geographical context: most of the stations located along the same river are found in the same group, with the only exception of the 4 stations along the Mississippi River. The further division of these 4 stations into 2 groups may be easily justified by the large geographical distance between the 2 groups: one group located in Minnesota (MN) and the other located in Illinois (IL).
In the supplementary material [5], regarding the real data demonstrations presented above, we have included more results with different choices of extremal subsample sizes, or say, top norm percentages. For the air pollution data, the results are relatively stable when the percentage varies. For the river discharge data, when the percentage is decreased to $5\%\sim 1\%$, the penalized ASW criterion starts to favor $k=5$. In this case, the station BELOW SPRAGUE RIVER NEAR CHILOQUIN labeled as group 2 in Table 1 is merged to the concomitant group labeled as 4 formed by SALEM, PORTLAND and HARRISBURG, all of which are geographically close (see Figure 10).
@@ 7 Summary and Discussion
Acknowledgment. We thank two anonymous reviewers for their helpful comments that lead to substantial improvements of the article.

### v4

@@ 6.2.2 River Discharge Data
The river discharge data concerns the daily discharge rate of rivers in North America sourced from the Global Runoff Data Centre [12]. The dataset comprises 16,386 daily records of discharge values from 13 stations spanning the period from December 1, 1976, to October 11, 2021. These 13 stations, shown in Table 1 and Fig. 10, are positioned along 5 rivers in America: Willamette River, Mississippi River, Williamson River, Hudson River, and Broad River.
Table 1: Clustering of 13 river discharge stations based on concomitant Extremes.
@@
Fig. 10: Geographical locations of the 13 river discharge stations.
Fig. 11: Penalized ASW curves for river discharge data (top $10\%$ norms).
Fig. 12: Estimated $B^{\top}$ for river discharge data (top $10\%$ norms).
As in the previous example, Fig. 12 presents the penalized ASW curves, from which we found that 6 seems to be an appropriate choice of order. Fig. 12 illustrates the factor matrix $B$ converted from the spectral estimation following the method in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$. In addition, for each row of the matrix $B$, we find to which factor index (the same as the cluster index in Fig. 12) the largest value corresponds. We include these factor indices in the last column of Table 1, which can be viewed roughly as markings of groups of concomitant extremes. The results align well with the expected geographical context: These 6 groups are in good accordance with the geographical context: most of the stations located along the same river are found in the same group, with the only exception of the 4 stations along the Mississippi River. The further division of these 4 stations into 2 groups may be easily justified by the large geographical distance between the 2 groups: one group located in Minnesota (MN) and the other located in Illinois (IL).
In the supplementary material, regarding the real data demonstrations presented above, we have included more results with different choices of extremal subsample sizes, or say, top norm percentages. For the air pollution data, the results are relatively stable when the percentage varies. For the river discharge data, when the percentage is decreased to $5\%\sim 1\%$, the penalized ASW criterion starts to favor $k=5$. In this case, the station BELOW SPRAGUE RIVER NEAR CHILOQUIN labeled as group 2 in Table 1 is merged to the concomitant group labeled as 4 formed by SALEM, PORTLAND and HARRISBURG, all of which are geographically close (see Fig. 10).
@@ 7 Summary and Discussion
Acknowledgment. We thank two anonymous reviewers and the associate editor for their helpful comments that lead to substantial improvements of the article.